CW1-CO3 DETERMINE THE THEVENINS AND NORTONS EQUIVALENT CIRCUITS OF THE FOLLOWING. 1. DETERMINE THE THEVENINS AND NORTONS
Answer #1

1 w 52 M А 2or Vab 2A lov 20 Calculation of va b= Uth: KCL at A v-10 + = 2 => 4V1 t = 4 => 1 + 4 => $V1 = 4 => Vi= 16 th=,= 1Z 1K2 m aku mm VA NI T 4 kr vth (1) 24A 4K 8 20v Vith:- Voltage at node A = V. KCL at A, ~ - 24+ 4 V-20+ š=0 => V-20+* + š =3 3) 1 oor w 15on m t I 24U zoon I=o. Vaba Vith Vith V200 = Voltage acron zoon 16V. = - 24x 200 200+100 abz Vth= (150 x0) +16